OpenStudy (anonymous):

m + 5 = 7.3 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5 A. m = –7.3 – 5 B. m = 5 – 7.3 C. m = 7.3 – 5

OpenStudy (anonymous):

The way you copied and paste this question is too confusing, you should attach a photo instead.

sammixboo (sammixboo):

Well to first solve this problem you multiply subtract 5 on both sides since you are adding 5 in the original equation. m + 5 = 7.3 m + 5 - 5 = 7.3 - 5 m = 7.3 - 5

OpenStudy (goldendragon01):

c